Suppose a glucose solution was prepared by dissolving 5.00 g of glucose in 100 g of water. At 363 K (90 degrees celcius). The experimentally determined vapor pressure of the glucose solution was found to be 69727.60 Pa. The vapor pressure of water at 363 K is 70100.90 Pa. Calculate the molar mass of glucose.



1
Expert's answer
2021-02-01T03:46:33-0500

number of moles of glucose = "\dfrac{\textsf{mass}}{\textsf{molar mass}}" = "\dfrac{5}{180}" = 0.028 moles


number of moles of water = "\dfrac{100}{18}" = 5.556 moles


Total number of moles of water = 5.556 + 0.028 = 5.584


molar fraction of solvent = "\dfrac{5.556}{5.584}" = 0.995


pressure fraction of solvent = "\dfrac{69727.60}{70100.90}" = 0.995.



This means the molar mass of glucose is truly 180 g/mol.